(4/216)^(-2/3) - (1/256)^(-3/4)
1
Expert's answer
2013-06-27T10:39:12-0400
(4/216)^(-2/3) - (1/256)^(-3/4)=(216/4)^(2/3) - (256)^(3/4)=54^(2/3) - (4^4)^(3/4)=54^(2/3) - 4^3=54^(2/3) - 64